# Integrable measure equivalence rigidity of right-angled Artin groups via quasi-isometry

### v1

@@ 2.2 Extension graphs and right-angled buildings
We emphasize that, unless $\Gamma$ is a complete graph, the extension graph $\Gamma^{e}$ is infinite and not locally finite. The conjugation action of $G_{\Gamma}$ on itself induces an action of $G_{\Gamma}$ on $\Gamma^{e}$ by graph automorphisms. Every standard flat $F\subseteq G_{\Gamma}$ determines a (finite) complete subgraph $\Delta(F)\subseteq\Gamma^{e}$, in the following way: $\mathrm{Stab}_{G_{\Gamma}}(F)$ is a parabolic subgroup, generated by a finite set of pairwise commuting infinite cyclic parabolic subgroups $\{Z_{1},\dots,Z_{k}\}$, and we let $\Delta(F)$ to the complete subgraph of $\Gamma^{e}$ spanned by the vertices corresponding to $Z_{1},\ldots,Z_{k}$. The number $k$ of vertices in $\Delta(F)$ is equal to the dimension of $F$. Given $\mathsf{v}\in V\Gamma^{e}$, a $\mathsf{v}$-line in $G_{\Gamma}$ is a standard line $\ell$ with $\Delta(\ell)=\{\mathsf{v}\}$. Note that a collection of pairwise distinct cyclic parabolic subgroups $P_{i}=g_{i}\langle v_{i}\rangle g_{i}^{-1}$ for $1\leq i\leq n$ mutually commute if and only $v_{i}$ and $v_{j}$ are adjacent in $\Gamma$ for $i\neq j$ and there exists $g\in G_{\Gamma}$ such that for every $i\in\{1,\dots,n\}$, one has $P_{i}=g\langle v_{i}\rangle g^{-1}$. This is a consequence of Proposition 2.1 and an induction on $n$. Thus, if $K$ is a complete subgraph of $\Gamma^{e}$, then there exists a standard flat $F=g\langle v_{1}\ldots,v_{n}\rangle$ such that $\Delta(F)=K$. If two standard flats $F_{1},F_{2}$ satisfy $\Delta(F_{1})=\Delta(F_{2})$, then they are parallel. Indeed, if $F_{i}=g_{i}G_{\Lambda_{i}}$ for every $i\in\{1,2\}$, then $\Delta(F_{1})=\Delta(F_{2})$ implies $g_{1}G_{\Lambda_{1}}g^{-1}_{1}=g_{2}G_{\Lambda_{2}}g^{-1}_{2}$. Then Proposition 2.1 implies that $F_{1}$ and $F_{2}$ are parallel. As a consequence, the map $\Delta$ from the collection of standard flats to the collection of cliques of $\Gamma^{e}$ induces a bijection between maximal standard flats in $G_{\Gamma}$ and maximal cliques in $\Gamma^{e}$. Thus we can define $\Gamma^{e}$ alternatively as follows. Vertices of $\Gamma^{e}$ are in one-to-one correspondence with parallelism classes of standard lines in $G_{\Gamma}$. Two vertices of $\Gamma^{e}$ are adjacent if and only if there are representatives of the associated parallelism classes that together span a $2$-dimensional standard flat.
@@
Recall that an interval in a partially ordered set $(\mathcal{P},\leq)$ is a subset of the form $I_{a,b}=\{x\in\mathcal{P}\mid a\leq x\leq b\}$ for some $a,b\in\mathcal{P}$ with $a\leq b$. If every interval in $\mathcal{P}$ is a Boolean lattice of finite rank, then there exists a unique (up to isomorphism) cube complex $|\mathcal{P}|$ whose poset of cubes is isomorphic to the poset of intervals of $\mathcal{P}$, see e.g. [AB08, Proposition A.38]. We call $|\mathcal{P}|$ the cubical realization of $\mathcal{P}$. We will be particularly interested in the case when $\mathcal{P}$ is the poset of standard flats in $G_{\Gamma}$, ordered by inclusion. Note that if we take $g_{1}G_{\Lambda_{1}}\leq g_{2}G_{\Lambda_{2}}$ in $\mathcal{P}$, then we can assume that $g_{2}=g_{1}$ up to changing coset representatives, and the interval between $g_{1}G_{\Lambda_{1}}$ and $g_{2}G_{\Lambda_{2}}$ consists of all cosets of form $g_{1}G_{\Lambda}$ with $\Lambda_{1}\subset\Lambda\subset\Lambda_{2}$. In particular it is a Boolean lattice of finite rank. The following notion was introduced by Davis in [Dav98].
@@
Let $\Gamma$ be a finite simplicial graph. The right-angled building of $G_{\Gamma}$, denoted by $\mathbb{B}_{\Gamma}$, is the cubical realization of the poset of standard flats in $G_{\Gamma}$.
@@
We comment on the terminology “right-angled building”, even though we will not explicitly need the connection to buildings in the rest of the paper. Davis [Dav98] explained that we can view the right-angled Artin group $G_{\Gamma}$ as a building modeled on a reflection group $W_{\Gamma}$ which is the right-angled Coxeter group with the same defining graph. Moreover, $\mathbb{B}_{\Gamma}$ is a geometric model witnessing some classical geometric properties of buildings, for example, any two points in $\mathbb{B}_{\Gamma}$ are contained in a common “apartment” which is a convex subcomplex isomorphic to the canonical $\mathrm{CAT}(0)$ cube complex (called the Davis complex) associated with $W_{\Gamma}$. We refer to [Dav98, Section 5] for a more detailed discussion.
By [Dav98, Corollary 11.7] (attributed to Meier), the cube complex $\mathbb{B}_{\Gamma}$ is $\mathrm{CAT}(0)$. There is a one-to-one correspondence between $k$-cubes in $\mathbb{B}_{\Gamma}$ and intervals of the form $I_{F_{1},F_{2}}$ where $F_{1}\subseteq F_{2}$ are two standard flats in $\Gamma$ with $\dim(F_{2})-\dim(F_{1})=k$. In particular, vertices of $\mathbb{B}_{\Gamma}$ correspond to standard flats in $\Gamma$. The rank of a vertex $v\in V\mathbb{B}_{\Gamma}$ is the dimension of the corresponding standard flat. We label every vertex of $\mathbb{B}_{\Gamma}$ by the type of the corresponding standard flat. The vertex set of $\mathbb{B}_{\Gamma}$ inherits a partial order from the poset of standard flats. There is an induced action of $G_{\Gamma}$ on $\mathbb{B}_{\Gamma}$ by cubical automorphisms, which preserves the labellings of vertices. This action is cocompact, but not proper – the stabilizer of a cube is isomorphic to $\mathbb{Z}^{n}$ where $n$ is the rank of the minimal vertex in this cube. Recall that the join of two simplicial graphs $\Gamma_{1}$ and $\Gamma_{2}$, denoted $\Gamma_{1}\circ\Gamma_{2}$, is the simplicial graph obtained from the disjoint union $\Gamma_{1}\sqcup\Gamma_{2}$ by adding an edge between any vertex in $\Gamma_{1}$ and any vertex in $\Gamma_{2}$. It follows from the definition that if $\Gamma=\Gamma_{1}\circ\Gamma_{2}$, then $\mathbb{B}_{\Gamma}\cong\mathbb{B}_{\Gamma_{1}}\times\mathbb{B}_{\Gamma_{2}}$.
@@
Let $x\in V\mathbb{B}_{\Gamma}$ be a vertex. Recall that the link of $x$ in $\mathbb{B}_{\Gamma}$, denoted by $\mathrm{lk}(x,\mathbb{B}_{\Gamma})$, is the simplicial complex formed by intersecting an $\varepsilon$-sphere around $x$ with $\mathbb{B}_{\Gamma}$, for $\varepsilon>0$ small enough. In particular, there is a bijection between vertices of $\mathbb{B}_{\Gamma}$ which are adjacent to $x$ and vertices in $\mathrm{lk}(x,\mathbb{B}_{\Gamma})$. Let $\mathrm{lk}^{+}(x,\mathbb{B}_{\Gamma})$ (resp. $\mathrm{lk}^{-}(x,\mathbb{B}_{\Gamma})$) be the full subcomplex of $\mathrm{lk}(x,\mathbb{B}_{\Gamma})$ spanned by vertices which are larger than $x$ (resp. smaller than $x$). Then $\mathrm{lk}(x,\mathbb{B}_{\Gamma})$ is the join of $\mathrm{lk}^{+}(x,\mathbb{B}_{\Gamma})$ and $\mathrm{lk}^{-}(x,\mathbb{B}_{\Gamma})$. Note that $\mathrm{lk}^{+}(x,\mathbb{B}_{\Gamma})$ is a finite complex. Suppose $x$ has rank $k$ and corresponds to a standard flat $F$ of dimension $k$. Vertices in $\mathrm{lk}^{-}(x,\mathbb{B}_{\Gamma})$ are in one-to-one correspondence with co-dimension 1 standard flats in $F$. Thus these vertices can be divided into $k$ different classes $V_{1},\dots,V_{k}$, corresponding to the $k$ parallelism classes of co-dimension 1 standard flats in $F$. Each $V_{i}$ is an infinite set. Moreover, $\mathrm{lk}^{-}(x,\mathbb{B}_{\Gamma})$ is a join of $k$ discrete sets $V_{1}*V_{2}*\cdots*V_{k}$. Thus two vertices of $\mathbb{B}_{\Gamma}$ with different ranks must have non-isomorphic links in $\mathbb{B}_{\Gamma}$. Now the lemma follows. ∎
@@ 2.3 Automorphisms of the extension graph and the right-angled building
We endow $\mathrm{Aut}(\mathbb{B}_{\Gamma})$ and $\mathrm{Aut}(\Gamma^{e})$ with the compact-open topology, making them Polish groups. A bijection $f:G_{\Gamma}\to G_{\Gamma}$ is flat-preserving if both $f$ and $f^{-1}$ send any standard flat bijectively onto another standard flat. Let $\mathrm{Bij}_{\operatorname{FP}}(G_{\Gamma})$ be the group of flat-preserving bijections of $G_{\Gamma}$, again equipped with the compact-open topology, or equivalently the topology of pointwise convergence, which makes it a Polish group. Lemma 2.5 implies that for every $f\in\mathrm{Aut}(\mathbb{B}_{\Gamma})$, the restriction of $f$ to the set of rank 00 vertices is a flat-preserving bijection of $G_{\Gamma}$. Conversely, any flat-preserving bijection of $G_{\Gamma}$ induces an automorphism of the poset of standard flats, hence an automorphism of $\mathbb{B}_{\Gamma}$. This yields an isomorphism of Polish groups $\mathrm{Aut}(\mathbb{B}_{\Gamma})\simeq\mathrm{Bij}_{\operatorname{FP}}(G_{\Gamma})$. We now explain how to identify $\mathrm{Aut}(\mathbb{B}_{\Gamma})$ to $\mathrm{Aut}(\Gamma^{e})$ when $|\operatorname{Out}(G_{\Gamma})|<+\infty$, as will be recorded in Lemma 2.6 below.
@@
We define a map $\Phi:\mathrm{Aut}(\mathbb{B}_{\Gamma})\to\mathrm{Aut}(\Gamma^{e})$ as follows. Take $\alpha\in\mathrm{Aut}(\mathbb{B}_{\Gamma})$. The restriction of $\alpha$ to the set of rank 00 vertices of $\mathbb{B}_{\Gamma}$ is a flat-preserving bijection $g:G_{\Gamma}\to G_{\Gamma}$. We claim that $g$ sends parallel standard lines to parallel standard lines. Indeed, let $\ell_{1}$ and $\ell_{2}$ be parallel standard lines. If $\ell_{1}$ and $\ell_{2}$ are contained in a common $2$-dimensional standard flat, then the claim is obvious. In general, it follows from Proposition 2.1 that there is a finite chain of standard lines starting from $\ell_{1}$ and ending at $\ell_{2}$ such that consecutive members in the chain are parallel and contained in a common 2-dimensional standard flat. The claim thus follows.

### v2

@@ 2.2 Extension graphs and right-angled buildings
We emphasize that, unless $\Gamma$ is a complete graph, the extension graph $\Gamma^{e}$ is infinite and not locally finite (assuming $\Gamma$ is connected). The conjugation action of $G_{\Gamma}$ on itself induces an action of $G_{\Gamma}$ on $\Gamma^{e}$ by graph automorphisms. Every standard flat $F\subseteq G_{\Gamma}$ determines a (finite) complete subgraph $\Delta(F)\subseteq\Gamma^{e}$, in the following way: $\mathrm{Stab}_{G_{\Gamma}}(F)$ is a parabolic subgroup, generated by a finite set of pairwise commuting infinite cyclic parabolic subgroups $\{Z_{1},\dots,Z_{k}\}$, and we let $\Delta(F)$ to the complete subgraph of $\Gamma^{e}$ spanned by the vertices corresponding to $Z_{1},\ldots,Z_{k}$. The number $k$ of vertices in $\Delta(F)$ is equal to the dimension of $F$. Given $\mathsf{v}\in V\Gamma^{e}$, a $\mathsf{v}$-line in $G_{\Gamma}$ is a standard line $\ell$ with $\Delta(\ell)=\{\mathsf{v}\}$.
Note that a collection of pairwise distinct cyclic parabolic subgroups $P_{i}=g_{i}\langle v_{i}\rangle g_{i}^{-1}$ for $1\leq i\leq n$ mutually commute if and only $v_{i}$ and $v_{j}$ are adjacent in $\Gamma$ for $i\neq j$ and there exists $g\in G_{\Gamma}$ such that for every $i\in\{1,\dots,n\}$, one has $P_{i}=g\langle v_{i}\rangle g^{-1}$. This is a consequence of Proposition 2.1 and an induction on $n$.
Thus, if $K$ is a complete subgraph of $\Gamma^{e}$, then there exists a standard flat $F=g\langle v_{1}\ldots,v_{n}\rangle$ such that $\Delta(F)=K$.
If two standard flats $F_{1},F_{2}$ satisfy $\Delta(F_{1})=\Delta(F_{2})$, then they are parallel. Indeed, if $F_{i}=g_{i}G_{\Lambda_{i}}$ for every $i\in\{1,2\}$, then $\Delta(F_{1})=\Delta(F_{2})$ implies $g_{1}G_{\Lambda_{1}}g^{-1}_{1}=g_{2}G_{\Lambda_{2}}g^{-1}_{2}$. Then Proposition 2.1 implies that $F_{1}$ and $F_{2}$ are parallel. As a consequence, the map $\Delta$ from the collection of standard flats to the collection of cliques of $\Gamma^{e}$ induces a bijection between maximal standard flats in $G_{\Gamma}$ and maximal cliques in $\Gamma^{e}$.
Thus we can define $\Gamma^{e}$ alternatively as follows. Vertices of $\Gamma^{e}$ are in one-to-one correspondence with parallelism classes of standard lines in $G_{\Gamma}$. Two vertices of $\Gamma^{e}$ are adjacent if and only if there are representatives of the associated parallelism classes that together span a $2$-dimensional standard flat.
@@
Recall that an interval in a partially ordered set $(\mathcal{P},\leq)$ is a subset of the form $I_{a,b}=\{x\in\mathcal{P}\mid a\leq x\leq b\}$ for some $a,b\in\mathcal{P}$ with $a\leq b$. If every interval in $\mathcal{P}$ is a Boolean lattice of finite rank, then there exists a unique (up to isomorphism) cube complex $|\mathcal{P}|$ whose poset of cubes is isomorphic to the poset of intervals of $\mathcal{P}$, see e.g. [AB08, Proposition A.38]. We call $|\mathcal{P}|$ the cubical realization of $\mathcal{P}$.
We will be particularly interested in the case when $\mathcal{P}$ is the poset of standard flats in $G_{\Gamma}$, ordered by inclusion. Note that if we take $g_{1}G_{\Lambda_{1}}\leq g_{2}G_{\Lambda_{2}}$ in $\mathcal{P}$, then we can assume that $g_{2}=g_{1}$ up to changing coset representatives, and the interval between $g_{1}G_{\Lambda_{1}}$ and $g_{2}G_{\Lambda_{2}}$ consists of all cosets of form $g_{1}G_{\Lambda}$ with $\Lambda_{1}\subset\Lambda\subset\Lambda_{2}$. In particular it is a Boolean lattice of finite rank. The following notion was introduced by Davis in [Dav98].
@@
Let $\Gamma$ be a finite simplicial graph. The right-angled building of $G_{\Gamma}$, denoted by $\mathbb{B}_{\Gamma}$, is the cubical realization of the poset of standard flats in $G_{\Gamma}$. The left action of $G_{\Gamma}$ on itself induces an action of $G_{\Gamma}$ on $\mathbb{B}_{\Gamma}$ by cubical automorphisms.
In Figure 1 below, we draw a subcomplex of $\mathbb{B}_{\Gamma}$ when $\Gamma$ is a pentagon with consecutive vertices denoted by $a,b,c,d,e$. The vertex in Figure 2.3 labeled by $ab$ corresponds to the identity left coset of the subgroup generated by $a$ and $b$. Similarly, we define other vertices in Figure 2.3. The central vertex corresponds to the identity element in $G_{\Gamma}$. The subcomplex of $\mathbb{B}_{\Gamma}$ displayed in Figure 2.3 is a fundamental domain for the action $G_{\Gamma}\curvearrowright\mathbb{B}_{\Gamma}$.
Figure 1: Fundamental domain of the pentagon building.
@@
We comment on the terminology “right-angled building”, even though we will not explicitly need the connection to buildings in the rest of the paper. Davis [Dav98] explained that we can view the right-angled Artin group $G_{\Gamma}$ as a building in the sense of Tits (see [Dav98, Section 3]) modeled on a reflection group $W_{\Gamma}$ which is the right-angled Coxeter group with the same defining graph. Moreover, $\mathbb{B}_{\Gamma}$ is a geometric model witnessing some classical geometric properties of buildings, for example, any two points in $\mathbb{B}_{\Gamma}$ are contained in a common “apartment” which is a convex subcomplex isomorphic to the canonical $\mathrm{CAT}(0)$ cube complex (called the Davis complex) associated with $W_{\Gamma}$. We refer to [Dav98, Section 5] for a more detailed discussion.
By [Dav98, Corollary 11.7] (attributed to Meier), the cube complex $\mathbb{B}_{\Gamma}$ is $\mathrm{CAT}(0)$. There is a one-to-one correspondence between $k$-cubes in $\mathbb{B}_{\Gamma}$ and intervals of the form $I_{F_{1},F_{2}}$ where $F_{1}\subseteq F_{2}$ are two standard flats in $\Gamma$ with $\dim(F_{2})-\dim(F_{1})=k$. In particular, vertices of $\mathbb{B}_{\Gamma}$ correspond to standard flats in $G_{\Gamma}$. The rank of a vertex $v\in V\mathbb{B}_{\Gamma}$ is the dimension of the corresponding standard flat. We label every vertex of $\mathbb{B}_{\Gamma}$ by the type of the corresponding standard flat. The vertex set of $\mathbb{B}_{\Gamma}$ inherits a partial order from the poset of standard flats.
The action of $G_{\Gamma}$ on $\mathbb{B}_{\Gamma}$ preserves the labellings of vertices. This action is cocompact, but not proper – the stabilizer of a cube is isomorphic to $\mathbb{Z}^{n}$ where $n$ is the rank of the minimal vertex in this cube.
Recall that the join of two simplicial graphs $\Gamma_{1}$ and $\Gamma_{2}$, denoted $\Gamma_{1}\circ\Gamma_{2}$, is the simplicial graph obtained from the disjoint union $\Gamma_{1}\sqcup\Gamma_{2}$ by adding an edge between any vertex in $\Gamma_{1}$ and any vertex in $\Gamma_{2}$. It follows from the definition that if $\Gamma=\Gamma_{1}\circ\Gamma_{2}$, then $\mathbb{B}_{\Gamma}\cong\mathbb{B}_{\Gamma_{1}}\times\mathbb{B}_{\Gamma_{2}}$.
@@
Let $x\in V\mathbb{B}_{\Gamma}$ be a vertex. Recall that the link of $x$ in $\mathbb{B}_{\Gamma}$, denoted by $\mathrm{lk}(x,\mathbb{B}_{\Gamma})$, is the simplicial complex formed by intersecting an $\varepsilon$-sphere around $x$ with $\mathbb{B}_{\Gamma}$, for $\varepsilon>0$ small enough. In particular, there is a bijection between vertices of $\mathbb{B}_{\Gamma}$ which are adjacent to $x$ and vertices in $\mathrm{lk}(x,\mathbb{B}_{\Gamma})$. We claim that any two vertices of $\mathbb{B}_{\Gamma}$ with different ranks must have non-isomorphic links in $\mathbb{B}_{\Gamma}$, which implies the lemma.
Now we prove the claim. Let $\mathrm{lk}^{+}(x,\mathbb{B}_{\Gamma})$ (resp. $\mathrm{lk}^{-}(x,\mathbb{B}_{\Gamma})$) be the full subcomplex of $\mathrm{lk}(x,\mathbb{B}_{\Gamma})$ spanned by vertices which are larger than $x$ (resp. smaller than $x$). A simplex in $\mathrm{lk}^{+}(x,\mathbb{B}_{\Gamma})$ corresponds to the corner of a cube $C^{+}\subset\mathbb{B}_{\Gamma}$ corresponding to an interval of form $[x,y]$ for some vertex $y\geq x$. A simplex in $\mathrm{lk}^{-}(x,\mathbb{B}_{\Gamma})$ corresponds to the corner of a cube $C^{-}\subset\mathbb{B}_{\Gamma}$ corresponding to an interval of form $[z,x]$ for some vertex $z\leq x$. Then $C^{+}$ and $C^{-}$ span a bigger cube in $\mathbb{B}_{\Gamma}$ isometric to $C^{+}\times C^{-}$ corresponding to the interval $[z,y]$. Thus $\mathrm{lk}(x,\mathbb{B}_{\Gamma})$ is the join of $\mathrm{lk}^{+}(x,\mathbb{B}_{\Gamma})$ and $\mathrm{lk}^{-}(x,\mathbb{B}_{\Gamma})$. Note that $\mathrm{lk}^{+}(x,\mathbb{B}_{\Gamma})$ is a finite complex. Suppose $x$ has rank $k$ and corresponds to a standard flat $F$ of dimension $k$. Vertices in $\mathrm{lk}^{-}(x,\mathbb{B}_{\Gamma})$ are in one-to-one correspondence with co-dimension 1 standard flats in $F$. Thus these vertices can be divided into $k$ different classes $V_{1},\dots,V_{k}$, corresponding to the $k$ parallelism classes of co-dimension 1 standard flats in $F$. Each $V_{i}$ is an infinite set. Moreover, $\mathrm{lk}^{-}(x,\mathbb{B}_{\Gamma})$ is a join of $k$ discrete sets $V_{1}*V_{2}*\cdots*V_{k}$. Thus the claim follows. ∎
@@ 2.3 Automorphisms of the extension graph and the right-angled building
We endow $\mathrm{Aut}(\mathbb{B}_{\Gamma})$ and $\mathrm{Aut}(\Gamma^{e})$ with the compact-open topology, making them Polish groups. A bijection $f:G_{\Gamma}\to G_{\Gamma}$ is flat-preserving if both $f$ and $f^{-1}$ send any standard flat bijectively onto another standard flat. Let $\mathrm{Bij}_{\operatorname{FP}}(G_{\Gamma})$ be the group of flat-preserving bijections of $G_{\Gamma}$, again equipped with the compact-open topology, or equivalently the topology of pointwise convergence, which makes it a Polish group. Lemma 2.5 implies that for every $f\in\mathrm{Aut}(\mathbb{B}_{\Gamma})$, the restriction of $f$ to the set of rank 0 vertices is a flat-preserving bijection of $G_{\Gamma}$. Conversely, any flat-preserving bijection of $G_{\Gamma}$ induces an automorphism of the poset of standard flats, hence an automorphism of $\mathbb{B}_{\Gamma}$. This yields an isomorphism of Polish groups $\mathrm{Aut}(\mathbb{B}_{\Gamma})\simeq\mathrm{Bij}_{\operatorname{FP}}(G_{\Gamma})$. We now explain how to identify $\mathrm{Aut}(\mathbb{B}_{\Gamma})$ to $\mathrm{Aut}(\Gamma^{e})$ when $|\operatorname{Out}(G_{\Gamma})|<+\infty$, as will be recorded in Lemma 2.6 below.
@@
We define a map $\Phi:\mathrm{Aut}(\mathbb{B}_{\Gamma})\to\mathrm{Aut}(\Gamma^{e})$ as follows. Take $\alpha\in\mathrm{Aut}(\mathbb{B}_{\Gamma})$. The restriction of $\alpha$ to the set of rank 0 vertices of $\mathbb{B}_{\Gamma}$ is a flat-preserving bijection $g:G_{\Gamma}\to G_{\Gamma}$. We claim that $g$ sends parallel standard lines to parallel standard lines. Indeed, let $\ell_{1}$ and $\ell_{2}$ be parallel standard lines. If $\ell_{1}$ and $\ell_{2}$ are contained in a common $2$-dimensional standard flat, then the claim is obvious. In general, it follows from Proposition 2.1 that there is a finite chain of standard lines starting from $\ell_{1}$ and ending at $\ell_{2}$ such that consecutive members in the chain are parallel and contained in a common 2-dimensional standard flat. The claim thus follows.